OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue). A. 4/3 B. 2/3 C. 7/9 D. 4/27

OpenStudy (anonymous):

Huh? You mean first we choose a red marble? And if we did, what difference does it make since we return it.

OpenStudy (anonymous):

thats wat i dont understand

OpenStudy (anonymous):

4/9 * 3/9 =4/27

OpenStudy (anonymous):

4/9 * 3/9 =4/27

OpenStudy (anonymous):

thank you u mind helping me with another one matricked